OpenStudy (anonymous):

Find the radius of the circle whose equation is x² + y² = 4. 2 4 16

OpenStudy (zale101):

\(\LARGE x^2+y^2=r^2\) Now, what is the radius of x² + y² = 4 ?

rvc (rvc):

That is the standard equation of circle now compare it :)

OpenStudy (zale101):

if \(r^2= Number\), then to figure out the radius alone, we take the square root of both sides if \(\Large \sqrt{r^2}= \sqrt{Number}\) if \(\Large r= \sqrt{Number}\)

OpenStudy (anonymous):

16

OpenStudy (zale101):

No.

OpenStudy (anonymous):

4 squared is 16.

OpenStudy (zale101):

it would be 16 if your equation was: x² + y² = 256 then the radius would be: \(\Large r=\sqrt{256}=16\)

rvc (rvc):

bring the equation in a form where all the variables and constants are in the square form

OpenStudy (anonymous):

I don't understand.

OpenStudy (anonymous):

Is it 4 then?

rvc (rvc):

\[x^2+y^2=(16)^2\] considering @Zale101 example

rvc (rvc):

when we compare the equation with \[\x^2+y^2=r^2\] r=16

OpenStudy (zale101):

Here are some other example: x² + y² = 256 radius is 16; because x² + y² = (16)^2 x² + y² = 81, radius is 9; becausex² + y² = 9^2 x² + y² = 64, radius is 8, because x² + y² = 8^2

rvc (rvc):

ur equation : \[x^2+y^2=(?)^2\]

OpenStudy (zale101):

x² + y² = 4, the radius for this is..?

OpenStudy (anonymous):

2

OpenStudy (zale101):

Bingo!

rvc (rvc):

All the best!!! :)

OpenStudy (anonymous):

Thanks!

OpenStudy (zale101):

x² + y² = 4, the radius is 2 because x² + y² = 2^2

OpenStudy (zale101):

Anytime

OpenStudy (anonymous):

wow too long :)